OpenStudy (anonymous):

Integral cos^2 x sin^4 x dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\cos^2xsin^4x dx\]

zepdrix (zepdrix):

Hmm they're both even powers, that's unfortunate. We'll have to do a lot of work to solve this one. We'll start by writing everything in terms of sine. So make this change first \(\Large \cos^2x \quad=\quad 1-\sin^2x\) and multiply it out. From there we can either apply `Half-Angle Formulas` a bunch of times. Or we can jump right into the `Sine Reduction Formula`. Have you learned about that yet?

zepdrix (zepdrix):

\[\large \int\limits \sin^n x\;dx \quad=\quad -\frac{1}{n}\sin^{n-1}x \cos x +\frac{n-1}{n}\int\limits \sin^{n-2}x\;dx\] Does this look familiar or no? :)

OpenStudy (anonymous):

Definately not

zepdrix (zepdrix):

Hmmm ok :( Well it would certainly be easier but if we have to do Half Angles, then so be it. :P

zepdrix (zepdrix):

\[\Large \int\limits \sin^4x-\sin^6x\;dx\]So ummm.

zepdrix (zepdrix):

The Sine Half-Angle Identity:\[\Large \sin^2x \quad=\quad\frac{1}{2}(1-\cos2x)\]

zepdrix (zepdrix):

So we can write sin^4 as,\[\Large \sin^4x\quad=\quad \left[\sin^2x\right]^2\quad=\quad \left[\frac{1}{2}(1-\cos2x)\right]^2\]

zepdrix (zepdrix):

Then multiply it all out,\[\Large \frac{1}{4}(1-2\cos2x+\cos^22x)\]

zepdrix (zepdrix):

Then apply the Cosine Half-Angle Identity to the last term there.\[\Large \frac{1}{4}(1-2\cos2x+\color{royalblue}{\cos^22x})\] \[\Large =\frac{1}{4}(1-2\cos2x+\color{royalblue}{\frac{1}{2}(1+\cos4x)})\]

zepdrix (zepdrix):

Now everything is written in terms of first powers. So they can be easily integrated, yay!

OpenStudy (anonymous):

Hahaha! Do you multiply out first?

zepdrix (zepdrix):

yes, distribute the fractions.\[\Large \int\limits \sin^4x-\sin^6x\;dx\] That takes care of the sin^4x portion. What to do about the sin^6x, hmmmm.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

The Equation is:\[\int\limits_{}^{}\cos^2xsin^4xdx\]

zepdrix (zepdrix):

Oh you -_- you and your ways!

zepdrix (zepdrix):

\[\large \int\limits \cos^2x \sin^4x\;dx\quad=\quad \int\limits (1-\sin^2x)\sin^4x\;dx \quad=\quad \int\limits \sin^4x-\sin^6x\;dx\]

OpenStudy (anonymous):

Ah! Ok =) Now what can we do with sin^6?

zepdrix (zepdrix):

hmmmm

zepdrix (zepdrix):

hmmmmmmmmm

OpenStudy (anonymous):

Is it possible to do like \[\int\limits_{}^{} (\sin^2xcosx)^2\] and apply a u-substitution?